# Any2Caption [Uncaptioned image]: Interpreting Any Condition to Caption for Controllable Video Generation

### v1

@@ 6.1 Setups
We manually construct 200 test cases for each type of condition (i.e., depth, human pose, multiple identities, camera, and compositional conditions) to evaluate the model’s performance. Additionally, we assess the model on publicly available benchmarks (e.g., [68, 41, 78]). For further details, please refer to the Appendix §$\S$§F.
@@
We leverage Qwen2VL-7B [56] as the backbone of our model, which supports both image and video understanding. The human pose in the input conditions is encoded and processed in the video format. The camera encoder adopts the vision encoder architecture with the following settings: the input channel of 96, the patch size of 16, the depth of 8, and 8 attention heads. During training, to simulate the brevity and randomness of user inputs, we randomly drop sentences from the short caption with a dropout rate of 0.6; a similar dropout strategy is applied to non-textual conditions. We conducted the training on 8$\times$×A800 GPUs. For further details on the training parameters for each stage, please refer to the Appendix §$\S$§F.
@@
Ctrl-Adapter [41] 20.37 - - - - - 25.63 - 94.53 20.73 4.63 46.98
@@ 6.2 Experimental Results and Analyses
To further showcase the model’s capacity to understand and leverage input conditions, we directly feed the structured captions—derived from our model’s interpretation—into downstream text-to-video generation systems (e.g., CogvideoX-2B [67] and Hunyuan [34]), as illustrated in Fig. 5. Even without explicit visual conditions (e.g., identities, depth, or camera movement), the resulting videos clearly align with the input prompts, such as the hat’s color and style, or the woman’s clothing texture in Example 1, indicating that our structured captions successfully capture intricate visual details. In particular, the model is able to accurately grasp dense conditions, such as depth sequences or compositional requirements in Example 4, ultimately enabling controllable video generation. While certain fine-grained elements may not be exhaustively described in text, resulting in occasional discrepancies with the actual visual content, the overall controllability remains robust.
@@
Here, we investigate whether integrating structured captions consistently improves controllable video generation in multiple methods. We explore the impact of adding camera, depth, identities, and human pose conditions to various controllable video generation methods. As shown in Tab. 5, all tested models exhibit consistent gains in overall video quality, such as smoothness and frame fidelity, after incorporating structured captions, without requiring any changes to the model architectures or additional training. Moreover, these models show enhanced adherence to the specified conditions, suggesting that our generated captions precisely capture user requirements and lead to more accurate, visually coherent video outputs. More examples can be found in the Appendix §$\S$§G.2.
@@
We examine whether a structured caption design is essential. We compare our structured caption approach with a simpler method, where we first caption the input condition (e.g., multiple identity images) and then concatenate that caption with the original short prompt, as shown in Tab. 3. Our results indicate that merely appending the condition’s caption to the short prompt can reduce video smoothness and image quality. One likely reason is that the identity images may contain extraneous details beyond the target subject, potentially conflicting with the original prompt and causing inconsistencies. Consequently, controllability in the final output is compromised. In contrast, our structured caption method accurately identifies the target subject and augments the prompt with relevant information, yielding more controllable video generation (cf. §$\S$§G.2).
Figure 6: Quantitative results on unseen conditions (i.e., segmentation [58], style [68], masked image [58], and sketch [58]) when using short and structured captions, respectively.
@@ 7 Conclusion
  • Limitation in Section §$\S$§A.
  • Ethic Statement results in Section §$\S$§B.
  • Extended Related Work in Section §$\S$§C.
  • Extended Dataset Construction Details in Section §$\S$§D.
  • More Statistics Information of IRScore in Section §$\S$§E.
  • Detailed Setups in Section §$\S$§F.
  • Extended Experiment Results and Analyses in Section §$\S$§G.
@@ C.1 Text-to-Video Generation
The development of video generation models has progressed from early GAN- and VAE-based approaches [7, 59, 20, 12, 19] to the increasingly popular diffusion-based methods [6, 5, 73, 49]. Among these, diffusion-in-transformer (DiT) architectures, such as OpenAI’s Sora [3] and HunyuanVideo [34], have demonstrated remarkable performance, producing photorealistic videos over extended durations. Controllable video generation [53, 11, 16, 21] has become an essential aspect of this field. Initially, research efforts centered predominantly on text-to-video generation [52, 24], which remains the most common approach. However, relying solely on text prompts can be insufficient for accurately capturing user intent, prompting exploration into other conditioning inputs such as static images [62, 17], user sketches [74, 58], human poses [77, 45, 33], camera perspectives [75, 21], and even additional videos [32, 14, 70]. Given this diversity of potential conditions, unifying them into a single “any-condition” video generation framework is highly valuable.
@@ C.2 Controllable Video Generation
Recent methods like VideoComposer [58], Ctrl-Adapter [41], and ControlVideo [74] have investigated any-condition video generation. Nevertheless, they still struggle with comprehensive controllability due to the complexity of multiple modalities and the limited capacity of standard diffusion or DiT encoders to interpret them. Inspired by the strong multimodal reasoning capabilities of modern MLLMs [43, 39, 56], we propose leveraging an MLLM to consolidate all possible conditions into a unified reasoning process, producing structured dense captions as inputs to a backbone Video DiT. SoTA DiT models already exhibit the capacity to interpret dense textual descriptions, as long as the input captions are sufficiently detailed in depicting both the scene and the intended generation goals. Building on this, our MLLM-based condition encoder directly addresses the comprehension bottleneck, theoretically enabling higher-quality video generation. To our knowledge, this work is the first to develop an MLLM specifically tailored for any-condition video generation. Because the caption-generation mechanism is decoupled from DiT, our proposed Any2Caption can be integrated into existing DiT-based methods without additional retraining.
@@ C.3 Video Captioning
Our approach is closely related to video recaptioning research, as our MLLM must produce dense captions based on the given conditions. In text-to-video settings, prior work [15, 67, 47, 27] has demonstrated the benefits of recaptioning videos to obtain more detailed textual annotations, thereby improving the training of longer and higher-quality video generation via DiT. ShareGPT4Video [10], for example, employs GPT-4V [1] to reinterpret video content and produce richer captions. MiraData [30] introduces structured dense recaptioning, while InstanceCap [15] focuses on instance-consistent dense recaptioning. Although we also pursue structured dense captions to enhance generation quality, our method diverges fundamentally from these previous approaches. First, because DiT models are already sufficiently powerful, we directly adopt an off-the-shelf Video DiT without incurring the substantial cost of training it with dense captions. Instead, we train an MLLM as an any-condition encoder at a comparatively lower cost; in the text-to-video scenario, for instance, we only need to train on pairs of short and dense captions, which are far easier and more abundant to obtain. Second, prior methods that recapturing the entire video risk introducing noise or even hallucinated content due to the current limitations of MLLMs in video understanding, potentially undermining DiT training quality, whereas our framework avoids this issue. Most importantly, while these approaches may rely on dense-caption-trained DiT models, real-world inference often involves very concise user prompts, creating a mismatch that can diminish final generation quality.
@@ C.4 Multimodal Large Language Models
Recent advances in Large Language Models (LLMs) [64] have catalyzed a surge of interest in extending their capabilities to multimodal domains [39, 37, 38, 61, 72]. A number of works integrate a vision encoder (e.g., CLIP [43], DINOv2[55], OpenCLIP[38]) with an LLM, often through a lightweight “connector” module (e.g., MLP [43], Q-former [35]), enabling the model to process image and video inputs with minimal additional training data and parameters [43, 63, 39]. These approaches have demonstrated promising performance on tasks such as image captioning, visual question answering, and video understanding. Beyond purely visual data, some researchers have investigated broader modalities, such as 3D motion [9, 28] or audio [13, 63], thereby expanding the application range of multimodal LLMs. Despite these advances, most existing MLLMs focus on a limited set of visual modalities and do not readily accommodate more specialized inputs like human pose or camera motion. This gap restricts their ability to handle more diverse and complex conditions in the controllable video generation field. In contrast, our work targets a broader spectrum of modalities, aiming to build a single model capable of interpreting and unifying image, video, human pose, and camera conditions. Specifically, we augment an existing MLLM with dedicated encoders for motion and camera features, thereby equipping the model to process arbitrary multimodal conditions and facilitate controllable video generation.
@@ F.1 Detailed Testing Dataset
Training Data Camera Movement Description Dataset (Manual) Action Description Dataset (Manual) MultiIDs Human Pose LLaVA-150K Camera LLaVA-150K Depth LLaVA-150K Alpaca-50K
@@ F.3 Detailed Implementations
In Tab. 11, we list the detailed hyperparameter settings in two stages. All the training is conducted on 8$\times$×A800 (80G) GPUs.
@@ G.2 More Video Visualization
Here, we compare the results of different customized models [31] after integrating structured captions. Fig. 14 shows camera-controlled video generation results, and Fig. 15 illustrates depth-controlled outcomes. We observe that structured captions improve image quality and motion smoothness by providing richer scene details.

### v2

@@ 6.1 Setups
[…]
@@
[…]
@@
Ctrl-Adapter [Ctrl-Adapter] 20.37 - - - - - 25.63 - 94.53 20.73 4.63 46.98
@@ 6.2 Experimental Results and Analyses
To further showcase the model’s capacity to understand and leverage input conditions, we directly feed the structured captions—derived from our model’s interpretation—into downstream text-to-video generation systems (e.g., CogvideoX-2B [CogVideoX] and Hunyuan [kong2024hunyuanvideo]), as illustrated in Fig. 5. Even without explicit visual conditions (e.g., identities, depth, or camera movement), the resulting videos clearly align with the input prompts, such as the hat’s color and style, or the woman’s clothing texture in Example 1, indicating that our structured captions successfully capture intricate visual details. In particular, the model is able to accurately grasp dense conditions, such as depth sequences or compositional requirements in Example 4, ultimately enabling controllable video generation. While certain fine-grained elements may not be exhaustively described in text, resulting in occasional discrepancies with the actual visual content, the overall controllability remains robust.
@@
[…]
@@
[…]
Figure 6: Quantitative results on unseen conditions (i.e., segmentation [VideoComposer-WangYZCWZSZZ23], style [StyleMaster], masked image [VideoComposer-WangYZCWZSZZ23], and sketch [VideoComposer-WangYZCWZSZZ23]) when using short and structured captions, respectively.
@@ 7 Conclusion
[…]
@@ C.1 Text-to-Video Generation
The development of video generation models has progressed from early GAN- and VAE-based approaches [BrooksHAWAL0EK22, WangBBD20, HaimFGSBDI22, ChuXMLT20, GurBW20] to the increasingly popular diffusion-based methods [BlattmannRLD0FK23, abs-2311-15127, ZhangXZFC24, QingZWWWZGS24]. Among these, diffusion-in-transformer (DiT) architectures, such as OpenAI’s Sora [sora] and HunyuanVideo [kong2024hunyuanvideo], have demonstrated remarkable performance, producing photorealistic videos over extended durations. Controllable video generation [abs-2411-04928, abs-2408-17424, abs-2411-18281, CameraCtrl] has become an essential aspect of this field. Initially, research efforts centered predominantly on text-to-video generation [Make-A-Video, CogVideo-Hong0ZL023], which remains the most common approach. However, relying solely on text prompts can be insufficient for accurately capturing user intent, prompting exploration into other conditioning inputs such as static images [Tune-A-Video-WuGWLGSHSQS23, I2V-Adapter-GuoZHGDWZLHZHM24], user sketches [ControlVideo, VideoComposer-WangYZCWZSZZ23], human poses [PoseCrafter-ZhongZYYZL24, MaHCWC0C24, DreamPose-KarrasHWK23], camera perspectives [CamI2V, CameraCtrl], and even additional videos [KaraKYRY24, DengWZTT24, I2VGen-XL]. Given this diversity of potential conditions, unifying them into a single “any-condition” video generation framework is highly valuable.
@@ C.2 Controllable Video Generation
Recent methods like VideoComposer [VideoComposer-WangYZCWZSZZ23], Ctrl-Adapter [Ctrl-Adapter], and ControlVideo [ControlVideo] have investigated any-condition video generation. Nevertheless, they still struggle with comprehensive controllability due to the complexity of multiple modalities and the limited capacity of standard diffusion or DiT encoders to interpret them. Inspired by the strong multimodal reasoning capabilities of modern MLLMs [llava-LiuLLL24, Video-LLaVA-LinYZCNJ024, Qwen2-VL], we propose leveraging an MLLM to consolidate all possible conditions into a unified reasoning process, producing structured dense captions as inputs to a backbone Video DiT. SoTA DiT models already exhibit the capacity to interpret dense textual descriptions, as long as the input captions are sufficiently detailed in depicting both the scene and the intended generation goals. Building on this, our MLLM-based condition encoder directly addresses the comprehension bottleneck, theoretically enabling higher-quality video generation. To our knowledge, this work is the first to develop an MLLM specifically tailored for any-condition video generation. Because the caption-generation mechanism is decoupled from DiT, our proposed Any2Caption can be integrated into existing DiT-based methods without additional retraining.
@@ C.3 Video Captioning
Our approach is closely related to video recaptioning research, as our MLLM must produce dense captions based on the given conditions. In text-to-video settings, prior work [InstanceCap, CogVideoX, abs-2407-02371, IslamHYNTB24] has demonstrated the benefits of recaptioning videos to obtain more detailed textual annotations, thereby improving the training of longer and higher-quality video generation via DiT. ShareGPT4Video [ShareGPT4Video], for example, employs GPT-4V [Gpt-4v] to reinterpret video content and produce richer captions. MiraData [MiraData-Ju00YWZX0S24] introduces structured dense recaptioning, while InstanceCap [InstanceCap] focuses on instance-consistent dense recaptioning. Although we also pursue structured dense captions to enhance generation quality, our method diverges fundamentally from these previous approaches. First, because DiT models are already sufficiently powerful, we directly adopt an off-the-shelf Video DiT without incurring the substantial cost of training it with dense captions. Instead, we train an MLLM as an any-condition encoder at a comparatively lower cost; in the text-to-video scenario, for instance, we only need to train on pairs of short and dense captions, which are far easier and more abundant to obtain. Second, prior methods that recapturing the entire video risk introducing noise or even hallucinated content due to the current limitations of MLLMs in video understanding, potentially undermining DiT training quality, whereas our framework avoids this issue. Most importantly, while these approaches may rely on dense-caption-trained DiT models, real-world inference often involves very concise user prompts, creating a mismatch that can diminish final generation quality.
@@ C.4 Multimodal Large Language Models
Recent advances in Large Language Models (LLMs) [Qwen2] have catalyzed a surge of interest in extending their capabilities to multimodal domains [Video-LLaVA-LinYZCNJ024, LLaMA-VID-LiWJ24, Mini-Gemini-abs-2403-18814, WuZXL00WZLL0QD24, OMG-LLaVA-ZhangL0YWJLY24]. A number of works integrate a vision encoder (e.g., CLIP [llava-LiuLLL24], DINOv2[Cambrian-1-TongBWWIAYYMWPF24], OpenCLIP[Mini-Gemini-abs-2403-18814]) with an LLM, often through a lightweight “connector” module (e.g., MLP [llava-LiuLLL24], Q-former [BLIP-2-0008LSH23]), enabling the model to process image and video inputs with minimal additional training data and parameters [llava-LiuLLL24, NExT-GPT-Wu0Q0C24, Video-LLaVA-LinYZCNJ024]. These approaches have demonstrated promising performance on tasks such as image captioning, visual question answering, and video understanding. Beyond purely visual data, some researchers have investigated broader modalities, such as 3D motion [MotionLLM-abs-2405-20340, MotionGPT-JiangCLYYC23] or audio [Qwen2-Audio, NExT-GPT-Wu0Q0C24], thereby expanding the application range of multimodal LLMs. Despite these advances, most existing MLLMs focus on a limited set of visual modalities and do not readily accommodate more specialized inputs like human pose or camera motion. This gap restricts their ability to handle more diverse and complex conditions in the controllable video generation field. In contrast, our work targets a broader spectrum of modalities, aiming to build a single model capable of interpreting and unifying image, video, human pose, and camera conditions. Specifically, we augment an existing MLLM with dedicated encoders for motion and camera features, thereby equipping the model to process arbitrary multimodal conditions and facilitate controllable video generation.
@@ F.1 Detailed Testing Dataset
Training Data
Camera Movement Description
Dataset (Manual)
Action Description
Dataset (Manual)
MultiIDs
Human Pose
LLaVA-150K
Camera
LLaVA-150K
Depth
LLaVA-150K
Alpaca-50K
@@ F.3 Detailed Implementations
In Tab. 11, we list the detailed hyperparameter settings in two stages. All the training is conducted on $\times8$0G GPUs.
@@ G.2 More Video Visualization
Here, we compare the results of different customized models [ju2025fulldit] after integrating structured captions. Fig. 14 shows camera-controlled video generation results, and Fig. 15 illustrates depth-controlled outcomes. We observe that structured captions improve image quality and motion smoothness by providing richer scene details.